OpenStudy (anonymous):

Evaluate: 5(6 − 3) − 1 • 6

OpenStudy (anonymous):

Is it 14?

OpenStudy (ajprincess):

is that 1*6 or 1.6

OpenStudy (anonymous):

15-6

OpenStudy (jackellyn):

Follow PEMDAS, Parenthesis, Exponents, Multiplication, Division, Addition Subtraction So, 5(3)-1*6 = 15-6 = 9

OpenStudy (anonymous):

Ok, Thanks. : )

OpenStudy (anonymous):

You should give me best answer lol